Please help me solve, a=?, b=35, c= 37
Answer by Theo(13342)

if this is a quadratic equation, then i believe -2 for a would do the trick.
you would get:
-2x^2 + 35x + 37 = 0 as the standard form of the quadratic equation.
factor this to get:
(-2x + 37) * (x + 1) = 0
when you multiply these factors together, you get:
-2x^2 - 2x + 37x + 37 = 0
when you combine like terms, you get:
-2x^2 + 35x + 37 = 0
that's about all i can think of in relation to this problem.
37 could only be 37 * 1
i didn't see any other combination that worked with that.
you then needed a to be equal to -2 in order to allow the addition of the a factors and the c factors to be added together to equal 35.
-2 + 37 = 35 so it looked liked an a factor of -1 would be good enough.
it turned out to be good.
this assumes i understand what you're trying to do correctly, which i believe is solve for a quadratic equation.